APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday November 25, 2002 (MJD 52604) ---=== OBSERVING TEAM ===--- Jurek Krzesinski Steph Snedden our swing (but he stayed working with us most of the night): Mike Harvanek all the shake- what else we can say: Scot Kleinman Eric Neilsen (by phone) ---=== OBSERVING PLAN ===--- Science plus night 2 of shakedown ---=== OBSERVING SUMMARY ===--- Test version of software declared current: iop v3_97_0, tpm v2_18_0, sdsscommish watcher v2_20_0 and interlocks v2_20_0, sdsshost watcher v2_20_0. Problems with the current version of interlocs on host (see the log). We finished Connie's bias/LED tests tonight and got one useful exposure on plate 677. The telescope, imager and spectro behaved properly. PRs closed: 4447 (SOP) gotoField has problems with gflat: ftclParseArg: Parameter type must be STRING. 3179 (MOP) autopicker should choose targets to meet a lenght of time rather than a number of targets 4361 (MOP) Autopicker runs out of targets too quickly with mtstds v3_5 PRs not closed: 3863 (SOP) efficiency meter too high (needs tests on clear sky) 4365 (MOP) ftelnetRead callers can get partial reads, and some callers (dfmRecv, at least) (needs more longer tests) 4385 (IOP) IOP murmur mnemonic table incomplete (assigned) ---=== OBSERVING LOG ===--- Afternoon checkout: Everything ok, but MIG values. Mike homed and relaxed mirrors and they went back to normal positions (exept E, which was a little off, so we are planing to home and relax them again if time permits). 00:55Z cart 6 plt 677 mounted. Again on slew we got a problem with gflat. Eric and Scot worked on this and made a fix. New IOP was cut. And now works properly. Also LoogTool misbehaves: we couldn't get MIG values with it - we are not sure what causes the problem. Eric told us that logTool and cleanupMurmur have to be started from within the same iop version. We tried this and it worked. Our observer account setup on hoggPT doesn't work anymore (logTool cannot read the software version on hoggPT) - it took Criag and Scot 1 hour to fix it. Stars on slew. Efficiency meter shows low values, but the weather deteriorates, so we'd better check it on a clear sky (PR 3863 not closed). However something is wrong - we've got too high S/N on that plate, as what we would have guessed from the efficiency meter. We need to check if the instrument block is read properly - and it looks like it is. We've got one useful exposure on this plate. We couldn't get S/N displayed on Plate Inventory Dtabase: Plan a Night's Observing page after updating S/N in the End of night page. See the problem section. PR field (4459). 02:12Z Clouds - an atmospheric front is coming - we might need to close the telescope soon. 02:44Z We put the enclosure over the telescope. At this point we started with Connie's checkBias checkLED. Bias taken in a dog house, tbars latched. LED drift, imagesr outside of the dog house, tbars unlatched. See results at the end of the night log. We declared current: On sdsscommish watcher v2_20_0 and interlocks v2_20_0 were declared current. On sdsshost, a ups list -a shows no interlocks v2_20_0. At the moment, watcher v2_20_0 is current, but there is no current interlocks. When you try to make a version current, you get the following: ups declare -c -f IRIX+6 interlocks v2_18_0 ERROR: No instance matches were made between the version file (/p/upsdb/interlocks/v2_20_0.version) and the table file (interlocks.table) for flavor (IRIX+6) and qualifiers () ERROR: Possible UPS database (/p/upsdb) corruption in product 'interlocks'. ERROR: No instance matches were made between the chain file (/p/upsdb/interlocks/test.chain) and the version file (v2_20_0.version) ERROR: Possible UPS database (/p/upsdb) corruption in product 'interlocks'. Que pasa? Also note, that what is in the software installs documentation for installing watcher and interlocks did not work as written. Instead of: commish> ups declare -c -f Linux+2 watcher tag_name commish> ups declare -c -f Linux+2 interlocks tag_name we had to use commish> ups declare -c -f Linux watcher tag_name commish> ups declare -c -f Linux interlocks tag_name Now, this may be because Robert checked things in with a flavor of "Linux", not "Linux+2", and normally the procedures will be able to be followed to the letter. Before we file a PR on this, we'd like feedback from Scot. iop v3_97_0, tpm v2_18_0 declared current. ---------------------------------------------- We also have now a tpm www pages: http://sdsshost.apo.nmsu.edu/tpm/ see for yourself. Connie's checkBias/checkLED --------------------------- (bias run: 3494, LET drift: 3495) im> checkBias 3494 30 /usrdevel/checkBias -doghouse -astro {65 WARN: } {65 FAIL: oddEvenAmp } {65 WARN: } {65 FAIL: oddEvenAmp } {71 WARN: leftBiasLevelComp rightBiasLevelComp } {71 FAIL: } {71 WARN: leftBiasLevelComp rightBiasLevelComp } {71 FAIL: } {15 WARN: rightRNComp } {15 FAIL: } {15 WARN: rightRNComp rightBiasVector} {15 FAIL: } {53 WARN: rightBiasVector} {53 FAIL: } {56 WARN: rightBiasVector} {56 FAIL: } {photo h5} {astro h6} checkLED -------- im> checkLED 3495 30 /usrdevel/checkBias {43 WARN: } {43 FAIL: leftGain } {46 WARN: leftGain } {46 FAIL: } Data are stored: sdss-commish: /skycat/imager/52604 --------------------------------------------------------------------- PRs closed: 4447 (SOP) gotoField has problems with gflat: ftclParseArg: Parameter type must be STRING. 3179 (MOP) autopicker should choose targets to meet a lenght of time rather than a number of targets. (tested) 4361 (MOP) Autopicker runs out of targets too quickly with mtstds v3_5 (tested) ----------------------------------------------------------------------- PRs not closed: 3863 (SOP) efficiency meter too high (needs tests on clear sky) 4365 (MOP) ftelnetRead callers can get partial reads, and some callers (dfmRecv, at least) (needs more longer tests) 4385 (IOP) IOP murmur mnemonic table incomplete (assigned) ------------------------------------------------------------------------ We left SOP PR 2561 to be checked after 10:00 am - Steph said, she will check it, when she'll recover from the night :). ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 3494 02:53Z 03:15Z 100 O -182.55 -176.97 51 ignore 3495 04:07Z 04:25Z 100 O -164.02 -159.46 43 calibration 3496 08:40Z 09:10Z 100 O -95.55 -88.01 57 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 05:27 17112 0.0 bias 05:29 17113 0.0 bias 05:47 17114 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:26Z 6 677 -18.75 0.0000 55.81 0.0000 164.23 0.0000 0.999910 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL 5751 Stays: JL 5752 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:25Z 6 677 0.99991 888 1849 -50 -18 55.9 0.8 9 268 gdr 1.5 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:37Z 32 16 10 263 (W) 511 - 01:07Z 33 15 9 265 (W) 371 - 01:34Z 33 15 7 265 (W) 334 - 02:05Z 33 15 9 266 (W) 308 - 02:38Z 33 15 8 266 (W) 331 - 03:10Z 33 15 7 265 (W) 331 - 03:43Z 33 15 8 285 (WNW) 289 - 04:15Z 33 17 7 266 (W) 304 - snow 04:47Z 33 17 9 264 (W) 321 - * * 05:19Z 32 17 9 183 (S) 398 - * 05:51Z 32 20 5 263 (W) 326 - * 06:23Z 30 21 0 157 (SSE) 372 - * 06:56Z 30 24 8 136 (SE) 337 - * 07:28Z 33 19 4 261 (W) 332 - * 08:00Z 33 21 2 202 (SSW) 324 - * 08:32Z 33 21 0 186 (S) 287 - * 09:04Z 31 27 14 215 (SW) 385 - * 09:34Z 30 28 6 170 (S) 283 - * 10:05Z 29 28 7 186 (S) 190 - * 10:35Z 30 29 9 244 (WSW) 12303 - * 11:05Z 29 29 6 260 (W) 7774 - * 11:36Z 29 29 2 216 (SW) 2206 - 11:59Z 29 29 1 242 (WSW) 16274 - ********* 12:30Z 29 29 2 184 (S) 47348 - 12:38Z 29 29 3 166 (SSE) 69705 - ---=== TELESCOPE STATUS ===--- 00:15Z telescope opened (fans were on an hour earlier). Dewars refilled. 02:44Z We put enclosure on the telescope, hoses hooked up. Fans off. ---=== SOFTWARE USED ===--- IOP/SOP: v3_96_0/v3_97_0 Watcher: v2_20_0 MCP: v5_16_0 TPM: tpm_v2_18_0 AstroDa: v14_47 TCC: 2.6.8 sdssProcedures: v1_56 SoS: v4_9_13 hoggPT: v1_5_8 plate-mapper: v4_1_0 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0890 0.0790 Axial B 0.7970 0.7970 Axial C 0.8080 0.8240 Trans D -9.0760 -9.1130 Lateral E 1.8390 1.8870 Lateral F 1.4060 1.4300 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5377. -3699. 923. -208. 31535. 30647. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: -1.3 Alt.: 30.018982 MIGS TONIGHT NOMINAL Axial A 1.5190 1.5170 Axial B 1.1640 1.1560 Axial C 1.1020 1.0990 Trans D -0.6960 -0.7070 GALILS Commanded: 1556651. 1492317. 1513331. -3400. -7950. Actual: 1619636. 1555273. 1576326. -3383. -7925. SETMIR VALUES SecDesOrient: 1257.00 0.00 -20.00 0.00 130.98 SecOrient: 1257.02 -0.03 -20.03 -0.49 130.90 ---=== PROBLEMS IN DETAIL ===--- Plate Inventory Dtabase: Plan a Night's Observing: -------------------------------------------------- When a (S/N)^2 value is entered in the end-of-night page of the plate database, and the page is updated, you see the (S/N)^2 proprly displayed on that page, BUT it is not displayed on the "plan a night's observing" page. Until this is fixed, we need to be very careful to check BOTH pages, so we don't take more exposures than we need.